# An Aligning and Training Framework for Multimodal Recommendations

### v1

@@ 6. Conclusion
A. Drawbacks of Current Multimodal Recommendation Methods
In this part, we analyze some drawbacks of current multimodal recommendation methods in detail. Following the discussions in Sec. 2, we select three typical baselines: VBPR, BM3 and FREEDOM.
VBPR is a representative of early methods that directly concatenate multimodal representations with ID representations for recommendation. However, these methods suffer from the semantic and distribution gap between different modalities. As illustrated in Fig. 5, when we recommend items based on the user interacted items using ID, image and text features, which are generated from a well-trained VBPR, the recommended items can be quite different. Recommending items with ID features will bring about target items that have similarities in usage. Recommending items with image features will result in target items having similar visual information. Recommending items with text features will make items with similar key phrases presented. Therefore, it is necessary to align and fuse all modality features for better recommendation.
BM3 is a representative of methods that train modality alignment losses together with the recommendation loss without utilizing our training strategies. We plot the training loss curves of the modality alignment loss (contrastive loss) and recommendation loss (BPR loss) in Figure 6(a). It shows that the recommendation loss quickly converged while the image/text contrastive loss decreased continuously over the training process. This is because multimodal features contain more complicated information, which requires more effort to learn. Simply optimizing these losses together like BM3 will make the recommendation loss converge before learning a good multimodal representation, leading to a relatively poor recommendation performance.
FREEDOM does not use alignment-related loss to incorporate vision and text information into learning ID embeddings. Instead, it tries to build similarity graphs for each modality feature and uses a weighted sum to acquire the multimodal similarity graph. It then uses this graph to aggregate features. Therefore, for each item, we calculate retrieved top-$K$ similar items based on vision similarity and text similarity separately, and get two retrieved item sets: $\mathcal{S}_{v},\mathcal{S}_{t}$. We then show the number of overlapped items in $\mathcal{S}_{v}$ and $\mathcal{S}_{t}$ in Figure 6(b) on Sports dataset. We discover that the overlapped items number is clearly lower than the fully overlapped number (the fully overlapped number is $K$ indeed), which indicates that items with similar text may not share similar images and they may conflict with each other. This modality incompatibility problem makes it hard for FREEDOM to align and fuse multimodal features with neural networks.
Figure 5. A case study of recommending items based on the user interacted item using ID, image or text features from VBPR.
(a) BM3 drawbacks.
(b) FREEDOM drawbacks.
Figure 6. Drawbacks of BM3 and FREEDOM.
B. Compared Baselines
@@ C. Hyper-parameter Settings
D. Analysis of Amazon Vision Features
In Sec. 5.3, we find that, when we mask 50% Amazon Baby vision features, the zero-shot recommendation results are better than those without masking any features (e.g., 0.0187 v.s. 0.0146 in R@50 in Table 3). To further investigate the reasons for this interesting phenomenon, we try to only use Amazon text features and ID features to train several baselines (i.e., do not use vision features), and compare the performance with their original results. Table  7 shows the detailed results on Baby dataset. We can observe that when we train baselines without Amazon vision features, the performance is still close to or even better than the original results. This is because the Amazon vision features contain less valuable information and do not contribute much to the final recommendation.
Baseline R@10 R@20 N@10 N@20
VBPR 0.0423 0.0663 0.0223 0.0284
VBPR w/o Vision 0.0458 0.0723 0.0242 0.0310
BM3 0.0564 0.0883 0.0301 0.0383
BM3 w/o Vision 0.0565 0.0880 0.0299 0.0385
FREEDOM 0.0624 0.0985 0.0324 0.0416
FREEDOM w/o Vision 0.0620 0.0993 0.0326 0.0422
E. Further Hyper-parameter Analysis
@@
In this part, we analyze the influence of the hyper-parameter $\lambda$, which controls the importance of the similarity-based regularizer, on Amazon Baby and Sports dataset. As illustrated in Fig. 7, we find that when this weight is very small, it cannot effectively control the similarity among content modalities, so the content modality representation will be optimized towards the ID-based representation without considering the content similarity, losing its own knowledge. Therefore, the performance is relatively bad. On the contrary, if $\lambda$ is large (e.g., $\lambda=1$), the whole optimization process will mainly focus on this task, ignoring other alignment tasks. As a result, the performance also decreases. In general, $\lambda$ reaches the best performance around 0.1.

### v4

@@ Acknowledgement
Details of Mask-Language-Modeling and Mask-Image-Modeling
We present the detailed loss formulation of mask-language-modeling (MLM) and mask-image-modeling (MIM), which are derived from the definitions in BERT (Devlin et al., 2018).
Suppose an item has text $t$ and image $v$. For $t$, we tokenize it and get $N$ text tokens $\mathcal{T}=\{t_{i}\}_{i=1}^{N}$ ($N$ is the pre-defined max length). Similarly, we split $v$ into $M$ image patches and tokenize it to $M$ visual tokens $\mathcal{V}=\{v_{j}\}_{j=1}^{M}$. We randomly mask approximately 15% text tokens and 40% image tokens, where the masked positions are denoted as $\mathcal{P}^{t}\in\{1,...,N\}^{0.15N}$ and $\mathcal{P}^{v}\in\{1,...,M\}^{0.4M}$, respectively. Next we replace the masked tokens with a unique text token $e^{t}$ and image token $e^{v}$, with learnable embeddings. Then the corrupted text tokens are $\hat{\mathcal{T}}=\{t_{i}|i\notin\mathcal{P}^{t}\}_{i=1}^{N}\cup\{e^{t}|i\in\mathcal{P}^{t}\}_{i=1}^{N}$; The corrupted image tokens are $\hat{\mathcal{V}}=\{v_{j}|j\notin\mathcal{P}^{v}\}_{j=1}^{M}\cup\{e^{v}|j\in\mathcal{P}^{v}\}_{j=1}^{M}$. The goal of inter-content alignment is to: (i) Predict the correct text tokens $z_{i}^{t}$ given corrupted text tokens $\hat{\mathcal{T}}$ and all image tokens $\mathcal{V}$, by minimizing $\mathcal{L}_{MLM}$; (ii) Predict the correct image tokens $z_{j}^{v}$ given corrupted image tokens $\hat{\mathcal{V}}$ and all text tokens $\mathcal{T}$, i.e., by minimizing $\mathcal{L}_{MIM}$. Formally, we use the maximum likelihood estimation to formulate them, where $p$ is an encoder with a softmax classifier:
$\mathcal{L}_{MLM}$ $=\max\mathbb{E}_{\mathcal{P}^{t}}\left[\sum_{i\in\mathcal{P}^{t}}\log p(z_{i}^{t}|\hat{\mathcal{T}}\cup\mathcal{V})\right]\,,$
$\mathcal{L}_{MIM}$ $=\max\mathbb{E}_{\mathcal{P}^{v}}\left[\sum_{j\in\mathcal{P}^{v}}\log p(z_{j}^{v}|\hat{\mathcal{V}}\cup\mathcal{T})\right]\,.$
Intuitively, $\mathcal{L}_{MLM}$ and $\mathcal{L}_{MIM}$ together aim to capture the relationship between image and text modality of an item. Then our MM encoder fuses the image and text representations to obtain the representation of the item.
Compared Baselines
@@
Table 8. Ablation study with mean and std values.
Dataset Variant R@10 R@20 N@10 N@20
Baby AlignRec 0.0673$\pm$ 0.0009 0.1039$\pm$0.0011 0.0365$\pm$0.0003 0.0458$\pm$0.0003
w/o CCA 0.0622$\pm$0.0008 0.0977$\pm$0.0011 0.0339$\pm$0.0005 0.0430$\pm$0.0006
w/o UIA 0.0639$\pm$0.0007 0.0996$\pm$0.0010 0.0345$\pm$0.0004 0.0440$\pm$0.0005
w/o REG 0.0603$\pm$0.0010 0.0940$\pm$0.0012 0.0321$\pm$0.0006 0.0412$\pm$0.0005
w/o Vision 0.0609$\pm$0.0009 0.0960$\pm$0.0013 0.0333$\pm$0.0003 0.0422$\pm$0.0006
w/o Text 0.0577$\pm$0.0005 0.0924$\pm$0.0010 0.0310$\pm$0.0003 0.0399$\pm$0.0005
Sports AlignRec 0.0764$\pm$0.0006 0.1158$\pm$0.0007 0.0416$\pm$0.0003 0.0517 $\pm$0.0003
w/o CCA 0.0664$\pm$0.0004 0.1028$\pm$0.0006 0.0355$\pm$0.0004 0.0449$\pm$0.0005
w/o UIA 0.0725$\pm$0.0003 0.1107$\pm$0.0005 0.0390$\pm$0.0005 0.0491$\pm$0.0005
w/o REG 0.0687$\pm$0.0005 0.1061$\pm$0.0006 0.0365$\pm$0.0003 0.0464$\pm$0.0004
w/o Vision 0.0744$\pm$0.0003 0.1122$\pm$0.0007 0.0399$\pm$0.0005 0.0496$\pm$0.0007
w/o Text 0.0675$\pm$0.0004 0.1024$\pm$0.0005 0.0369$\pm$0.0004 0.0460$\pm$0.0004
Hyper-parameter Settings
[…]
Analysis on Parameter Amount and Inference Speed
Due to the huge number of items and users, the structure’s complexity in multimodal recommendation can significantly influence its feasibility to deployment. Therefore, we present the parameter amount and inference speed of AlignRec and three well-performed baselines in Table 9. The inference time in this table means the total inference time over the test set. We can observe that the parameter amount of AlignRec is smaller than current baselines, with better performance. This is because our framework mainly focuses on the alignment design, training and evaluation, with small and simple network structures. Our multimodal representations are also smaller and more effective than before.
Table 9. Results on parameter amount and inference speed on four methods.
Dataset Method # Parameter Inference Time
Baby AlignRec 7.17M 2.15s
BM3 31.57M 2.32s
MGCN 35.58M 2.40s
FREEDOM 33.56M 2.22s
Sports AlignRec 17.60M 4.04s
BM3 82.98M 4.22s
MGCN 87.00M 4.26s
MGCN 85.96M 4.13s
Mean and STD Values in Ablation Study
To further validate the effectiveness of our framework, we conduct experiments on the ablation study with 5 seeds, and show the mean and std values in Table 8. We can observe that each component contributes to AlignRec.
Further Hyper-parameter Analysis
@@
Importance weight $\lambda$
It controls the importance of the similarity-based regularizer, on Amazon Baby and Sports dataset. As illustrated in Fig. 6, we find that when this weight is very small, it cannot effectively control the similarity among content modalities, so the content modality representation will be optimized towards the ID-based representation without considering the content similarity, losing its own knowledge. Therefore, the performance is relatively bad. On the contrary, if $\lambda$ is large (e.g., $\lambda=1$), the whole optimization process will mainly focus on this task, ignoring other alignment tasks. As a result, the performance also decreases. In general, $\lambda$ reaches the best performance around 0.1.
Number of layers $l$ in multimodal encoder
It determines the number of transformer layers in the multimodal encoder module, and is related to the scaling law of the multimodal recommendation. We change $l$ in $\{12,16,20\}$ ($l=12$ is the default setting in AlignRec), and the results are presented in Table 10. We can observe that when $l$ becomes larger, the performance slightly decreases. A larger $l$ represents more complicated networks in our multimodal encoder, which requires more data and more training time.
Table 10. Hyper-parameter study on the number of layers $l$ in the multimodal encoder.
Dataset $l$ R@10 R@20 N@10 N@20
Baby 12 0.0674 0.1046 0.0363 0.0458
16 0.0679 0.1052 0.0370 0.0463
20 0.0652 0.1027 0.0349 0.0445
Sports 12 0.0758 0.1160 0.0414 0.0517
16 0.0750 0.1154 0.0409 0.0512
20 0.0732 0.1129 0.0401 0.0503